A 1.0 dm3 glass bulb contains 1.0 × 1023 H2 molecules. If the pressure
exerted by the gas is 100 kPa, what are (a) the temperature of the gas, (b) the
root mean square speeds of the molecules? (c) Would the temperature be
different if they were O2 molecules?
1
Expert's answer
2020-10-26T14:47:53-0400

(a) n(H2) = 1*1023/6.02*1023 = 0.166 mol

p*V=nRT

T=p*V/(nR)

T = 100000*0.001/(0.166*8.314) = 72.5 K

(b) Vrms = (3*k*T/M)0.5 = (3*1.38*10-23*72.5/2)0.5 = 3.87*10-11 m

(c) yes, oxygen is a liquid under this preasure and temperature